_120) af terminal a to b al terminal and R These in theorem to find ob [20] Problem 4
Answer #1

VAh Vth Vth 4 1.57 from Cirluit A pply mesh aalysis ot ode tor loop i, 4li,-1:5 In) -6+ 3i, +4i, = 4li, - 1:si,) +ai= 6 -2i,+Isc.! Isc 1.5I, its neglecte d Short cirluit paratlel to So Isc 1151 from lir Cuit Iq= Isc from loop il i;=15 Isc mes h anal-2 Isc +3 Isc =6 Isc = 6 A Vth Rth Isc Rth 4.8 Rth Thwenins Equivalent carlit Rth =0-8 2 Vth=4.8V